# Hierarchical Prompting Taxonomy: A Universal Evaluation Framework for Large Language Models Aligned with Human Cognitive Principles

### v4

@@ Abstract
Assessing the effectiveness of large language models (LLMs) in performing different tasks is crucial for understanding their strengths and weaknesses. This paper presents Hierarchical Prompting Taxonomy (HPT), grounded on human cognitive principles and designed to assess LLMs by examining the cognitive demands of various tasks. The HPT utilizes the Hierarchical Prompting Framework (HPF), which structures five unique prompting strategies in a hierarchical order based on their cognitive requirement on LLMs when compared to human mental capabilities. It assesses the complexity of tasks with the Hierarchical Prompting Index (HPI), which demonstrates the cognitive competencies of LLMs across diverse datasets and offers insights into the cognitive demands that datasets place on different LLMs. This approach enables a comprehensive evaluation of an LLM’s problem-solving abilities and the intricacy of a dataset, offering a standardized metric for task complexity. Extensive experiments with multiple datasets and LLMs show that HPF enhances LLM performance by 2%-63% compared to baseline performance, with GSM8k being the most cognitively complex task among reasoning and coding tasks with an average HPI of 3.20 confirming the effectiveness of HPT. To support future research and reproducibility in this domain, the implementations of HPT and HPF are available here.
Codehttps://github.com/devichand579/HPT
1 Introduction
@@
Large Language Models (LLMs) have revolutionized natural language processing (NLP), enabling significant advancements in a wide range of applications. Conventional evaluation frameworks often apply a standard prompting approach to assess different LLMs, regardless of the complexity of the task, which may result in biased and suboptimal outcomes. Moreover, applying the same prompting approach across all samples within a dataset without considering each sample’s relative complexity adds to the unfair situation. To achieve a more balanced evaluation framework, it is essential to account for both the task-solving ability of LLMs and the varying cognitive complexities of the dataset samples. This limitation highlights the need for more sophisticated evaluation methods that can adapt to varying levels of task complexity.
@@
Figure 1: Hierarchical Prompting Framework includes five distinct prompting strategies, each designed for different levels of task complexity to ensure the appropriate prompt is selected for the given task. A $\checkmark$ indicates task completion, while a $\times$ signifies task incompletion.
@@
Within this study, complexity is defined as the cognitive demands associated with solving a task or the cognitive load introduced by a prompting strategy on LLMs. Henceforth, the term ”complexity” will be applied solely in this context. Task complexity, within the realm of human cognition, pertains to the cognitive requirements that a task imposes, which includes the diverse levels of mental effort necessary for processing, analyzing, and synthesizing information. According to Sweller (1988), tasks become more complex as they require greater cognitive resources, engaging working memory in more demanding processes such as reasoning and problem-solving. Similarly, Anderson et al. (2014) highlights that human cognitive abilities span a continuum from basic recall to higher-order thinking, with increasing difficulty correlating to tasks that demand analysis, synthesis, and evaluation. When applied to LLMs, the complexity of prompting strategies can be systematically evaluated by mapping them onto this human cognitive hierarchy. This alignment allows for an assessment of how LLMs perform tasks that reflect varying degrees of cognitive load, thereby providing a structured framework for understanding the cognitive demands associated with various tasks. By imposing this cognitive complexity framework on LLMs, this paper establishes a universal evaluation method, grounded in human cognitive principles, that enables more precise comparisons of model performance across tasks with varying levels of difficulty.
@@
This paper introduces the HPT, a set of rules that maps the human cognitive principles for assessing the complexity of different prompting strategies. It employs the HPF shown in Figure 1, a prompt selection framework that selects the prompt with the optimal cognitive load on LLM required in solving the task. HPF enhances the interaction with LLMs, and improves performance across various tasks by ensuring prompts resonate with human cognitive principles. The main contributions of this paper are as follows:
  1. 1.
@@
    Hierarchical Prompting Taxonomy (HPT): The paper proposes HPT, a set of rules that maps prompting strategies onto human cognitive principles, thereby enabling a universal measure of task complexity for LLMs.
  2. 2.
    Hierarchical Prompting Framework (HPF): The HPF is a framework designed to select the most effective prompt from five distinct prompting strategies, which optimizes the cognitive load on LLMs during task completion. This framework not only facilitates a more accurate evaluation of LLMs but also enhances their performance, delivering more transparent insights.
  3. 3.
@@
    Hierarchical Prompting Index (HPI): HPI [^1] HPI can be quantitatively assessed to analyze the cognitive abilities of an LLM and the cognitive demands imposed by datasets on LLMs, as both factors are interchangeably related to the complexity of tasks. quantitatively evaluates the task complexity of LLMs across various datasets, offering insights into the cognitive demands that each task imposes on different LLMs.
HPF can be effectively compared to an ”Open Book” examination as shown in Figure 2, where questions represent tasks and textbooks serve as prompting strategies. In this analogy, the exam questions vary in complexity, from simple factual recall to intricate analytical problems, analogous to the tasks in HPT that are evaluated based on their cognitive demands. Similarly, textbooks provide structured guidance for solving these questions, much like the HPF organizes prompts in increasing levels of complexity to support LLMs. For example, a straightforward glossary lookup corresponds to a low-complexity task, while solving a multi-step analytical problem requiring synthesis of concepts represents a high-complexity task. The effort a student invests in answering a question mirrors the HPI, which measures the cognitive load placed on the LLM. Just as students perform better with structured resources like textbooks, LLMs improve with well-designed hierarchical prompting strategies, enabling them to tackle progressively complex tasks effectively.
@@
Figure 2: Analogical framework comparing the HPF with ”Open Book” examination methodology. The diagram illustrates how HPF components (below) mirror traditional educational assessment elements (above), with parallel relationships between task complexity levels, resource utilization (prompts/textbooks), and performance metrics (HPI/student effort). This comparison demonstrates how LLM task complexity scales similarly to educational assessment complexity, from simple lookup tasks to complex synthesis problems
@@
The remainder of the paper is structured as follows: Section 2 reviews the related work on prompting and evaluation in LLMs. Section 3 details the HPT and its associated frameworks. Section 4 outlines the experimental setup, results, and ablation studies. Section 5 concludes the paper. Section 6 discusses the ethical impact of the work.
2 Related Work
@@ 2.1 Prompt Engineering
Prompting plays a vital role in unlocking the full potential of LLMs. By designing specific input prompts, the LLM’s responses can be guided, significantly influencing the quality and relevance of the output. Effective prompting strategies have enhanced LLM performance on tasks ranging from simple question-answering to complex reasoning and problem-solving. Recent research has explored various approaches to prompting and reasoning evaluation in LLMs. Chain-of-Thought (CoT) prompting (Wei et al. 2022b) elicits step-by-step reasoning, improving performance on complex tasks. Specializing smaller models (Fu et al. 2023) and using large models as reasoning teachers (Ho, Schmid, and Yun 2022) have demonstrated the potential for enhancing reasoning capabilities. Emergent abilities in LLMs, which appear suddenly at certain scale thresholds, have also been a topic of interest. Wei et al. (2022a) examined these abilities in few-shot prompting, discussing the underlying factors and implications for future scaling. Complementing this, Kojima et al. (2022) demonstrated that LLMs could exhibit multi-step reasoning capabilities in a zero-shot setting by simply modifying the prompt structure, thus highlighting their potential as general reasoning engines. Yao et al. (2023) introduced the Tree-of-Thoughts framework, enabling LLMs to deliberate over coherent text units and perform heuristic searches for complex reasoning tasks. This approach generalizes over chain-of-thought prompting and has shown significant performance improvements in tasks requiring planning and search, such as creative writing and problem-solving games. Kong et al. (2024) introduced role-play prompting to improve zero-shot reasoning by constructing role-immersion interactions, which implicitly trigger chain-of-thought processes and enhance performance across diverse reasoning benchmarks. Progressive-hint prompting (Zheng et al. 2023) has been proposed to conceptualize answer generation and guide LLMs toward correct responses. Metacognitive prompting (Wang and Zhao 2024) incorporates self-aware evaluations to enhance understanding abilities.
These works collectively highlight the advancements in leveraging LLMs through innovative prompting techniques, addressing their emergent abilities, reasoning capabilities, interaction strategies, robustness, and evaluation methodologies. Despite significant advancements, the current LLM research reveals several limitations, particularly in terms of prompt design, handling complex reasoning tasks, and evaluating model performance across diverse scenarios. While promising, the emergent abilities of LLMs often lack predictability and control, and the robustness of these LLMs in the face of misleading prompts remains a concern.
2.2 Prompt Optimization and Selection
@@
The challenge of optimizing prompts for LLMs has been addressed in several key studies, each contributing unique methodologies to enhance model performance and efficiency. Shen et al. (2023) introduce PFLAT, a metric utilizing flatness regularization to quantify prompt utility, which leads to improved results in classification tasks. Do et al. (2024) propose a structured three-step methodology that contains data clustering, prompt generation, and evaluation, effectively balancing generality and specificity in prompt selection. ProTeGi (Pryzant et al. 2023) offers a non-parametric approach inspired by gradient descent, leveraging natural language ”gradients” to iteratively refine prompts. Wang et al. (2024) present PromISe, which transforms prompt optimization into an explicit chain of thought, employing self-introspection and refinement techniques. Zhou et al. (2023b) proposed DYNAICL, a framework for efficient prompting that dynamically allocates in-context examples based on a meta-controller’s predictions, achieving better performance-efficiency trade-offs compared to uniform example allocation.
@@
These studies aim to automate prompt design, moving away from traditional manual trial-and-error methods while emphasizing efficiency and scalability across various tasks and models. They report significant improvements in LLMs performance, with enhancements ranging from 5% to 31% across different benchmarks. This body of work underscores the increasing importance of prompt optimization and selection in unlocking the potential of LLMs and points toward future research avenues, such as exploring theoretical foundations, integrating multiple optimization techniques, and distinguishing between task-specific and general-purpose strategies.
2.3 Evaluation Benchmarks
@@
To facilitate the evaluation and understanding of LLM capabilities, Zhu et al. (2024) introduced PromptBench, a unified library encompassing a variety of LLMs, datasets, evaluation protocols, and adversarial prompt attacks. This modular and extensible tool aims to support collaborative research and advance the comprehension of LLM strengths and weaknesses. Further exploring reasoning capabilities, Qiao et al. (2023) categorized various prompting methods and evaluated their effectiveness across different model scales and reasoning tasks, identifying key open questions for achieving robust and generalizable reasoning. (Wang et al. 2021) introduced a multi-task Benchmark for robustness Evaluation of LLMs extends the original GLUE (Wang et al. 2018) benchmark to assess model robustness against adversarial inputs. It incorporates perturbed versions of existing GLUE tasks, such as paraphrasing, negation, and noise, to test models’ abilities with challenging data. The study highlights that despite their success on clean datasets, state-of-the-art models often struggle with adversarial examples, underscoring the importance of robustness evaluations in model development.
3 Hierarchical Prompting Taxonomy
@@
Figure 3: Hierarchical Prompting Taxonomy: A taxonomy designed to assess the complexity of prompting strategies based on the criteria: Basic Recall and Reproduction, Understanding and Interpretation, Analysis and Reasoning, and Application of Knowledge and Reasoning.
3.1 Governing Rules
@@
  1. 1.
    Basic Recall and Reproduction: This reflects the fundamental cognitive process of remembering and reproducing factual information without analysis or interpretation, which involves mere recognition or retrieval of knowledge from memory (Anderson et al. 2014).
  2. 2.
    Understanding and Interpretation: This corresponds to the second cognitive rule of (Bloom 1956), where individuals must not only recall information but also explain it in their own words, summarize key points or clarify the meaning of content. This rule demands an intermediate cognitive load involving information processing rather than retrieving it.
  3. 3.
    Analysis and Reasoning: This aligns with the analysis stage of (Bloom 1956), which involves higher cognitive functions such as comparison, contrast, and deep understanding of the underlying principles. It is more complex than mere understanding because it requires examining structure and identifying patterns and connections.
  4. 4.
    Application of Knowledge and Execution: This mirrors the application and evaluation stages of (Bloom 1956), where individuals must not only understand and analyze but also use knowledge to perform multi-step tasks, solve complex problems, and execute decisions. It represents the most cognitively complex tasks, which require synthesis of information and practical decision-making, highlighting the critical leap from understanding theory to executing it in practice.
@@
3.2 Hierarchical Prompting Framework
@@
The set of five prompting strategies were chosen from a diverse range of existing strategies to populate the framework, guided by a human judgment policy, prioritizing comprehensiveness in cognitive demands rather than the sheer number of strategies. See Appendix A for more details. Consequently, the HPF can be replicated or expanded with other relevant prompting strategies that exhibit similar cognitive demands, making the framework adaptable. The five prompting strategies, listed from least to most complex, are as follows:
  1. 1.
    Role Prompting (Kong et al. 2024): Prompts that specify the role of the LLM in task resolution represent the lowest level of complexity within the framework, as they exhibit minimal influence from the four principles of the HPT.
  2. 2.
    Zero-Shot Chain-of-Thought Prompting (Zero-CoT) (Kojima et al. 2022): Prompts that use the phrase “Let’s think step by step” without providing prior examples aim to encourage critical thinking and problem-solving. This strategy implicitly activates chain-of-thought reasoning, demonstrating moderate influence from rule 3 of the HPT, while exerting a weaker influence from the other rules.
  3. 3.
    Three-Shot Chain-of-Thought Prompting (3-CoT) (Wei et al. 2022b): Prompts that present three examples to guide the LLM’s reasoning process exert a strong influence from rules 1 and 2, requiring recall and interpretation of the examples to complete the task. This strategy also demonstrates moderate influence from rule 3, increasing its cognitive demands.
  4. 4.
    Least-to-Most Prompting (Zhou et al. 2023a): A sequence of prompts that breaks the task into sub-problems exerts significant influence from all three rules 1, 2, and 3. This approach requires recalling prior prompts, interpreting previous responses, and analyzing them to effectively solve the task, resulting in a highly cognitively demanding strategy.
  5. 5.
    Generated Knowledge Prompting (GKP) (Liu et al. 2022): Prompts that require integrating external knowledge to generate relevant information represent the most complex and cognitively demanding strategy. This approach is strongly influenced by rules 2, 3, and 4, as it involves correlating knowledge with the prompt and applying and analyzing external information, making it the most cognitively demanding within the HPT framework. In the experiments, Llama-3 8B is used to generate external knowledge.
3.3 Hierarchical Prompting Index
HPI is an evaluation metric for assessing the task complexity of LLMs over different datasets, which is influenced by the HPT rules. A lower HPI for a dataset suggests that the corresponding LLM is more adept at solving the task with fewer cognition processes. For each dataset instance, we begin with the least complex prompting strategy and progressively move through the HPF prompting strategies until the instance is resolved. The HPI corresponds to the complexity level of the prompting strategy where the LLM first tackles the instance.
Algorithm 1 HPI Metric

### v5

@@ Abstract.
Assessing the effectiveness of large language models (LLMs) in performing different tasks is crucial for understanding their strengths and weaknesses. This paper presents Hierarchical Prompting Taxonomy (HPT), grounded on human cognitive principles and designed to assess LLMs by examining the cognitive demands of various tasks. The HPT utilizes the Hierarchical Prompting Framework (HPF), which structures five unique prompting strategies in a hierarchical order based on their cognitive requirement on LLMs when compared to human mental capabilities. It assesses the complexity of tasks with the Hierarchical Prompting Index (HPI), which demonstrates the cognitive competencies of LLMs across diverse datasets and offers insights into the cognitive demands that datasets place on different LLMs. This approach enables a comprehensive evaluation of LLM’s problem-solving abilities and the intricacy of a dataset, offering a standardized metric for task complexity. Extensive experiments with multiple datasets and LLMs show that HPF enhances LLM performance by $2\to 63\%$ compared to baseline performance, with GSM8k being the most cognitively complex task among reasoning and coding tasks with an average HPI of $3.20$ confirming the effectiveness of HPT. To support future research in this domain, the implementations of HPT and HPF are publicly available [^1] Code and Experiments.
Prompting Taxonomy, Cognitive Demands, Prompt Optimization
isbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Natural language generationccs: Computing methodologies Reasoning about belief and knowledgeccs: Computing methodologies Cognitive science
1. Introduction
@@
Large Language Models (LLMs) have revolutionized natural language processing (NLP), enabling significant advancements in a wide range of applications. Conventional evaluation frameworks often apply a standard prompting approach to assess different LLMs, regardless of the complexity of the task, which may result in biased and suboptimal outcomes. Moreover, applying the same prompting approach across all samples within a dataset without considering each sample’s relative complexity adds to the unfair situation. To achieve a more balanced evaluation framework, it is essential to account for both the task-solving ability of LLMs and the varying cognitive complexities of the dataset samples. This limitation highlights the need for more sophisticated evaluation methods that can adapt to varying levels of sample task complexity.
[…]
Hierarchical Prompting Framework diagram.
@@
Figure 1. The Hierarchical Prompting Framework includes five distinct prompting strategies, each designed for different levels of task complexity to ensure the appropriate prompt is selected for the given task. A $\checkmark$ indicates task completion, while a $\times$ signifies task incompletion.
@@
This study defines complexity as the cognitive demands imposed by a task or the cognitive load introduced by a prompting strategy on LLMs. Task complexity in human cognition reflects the mental effort required for processing, analyzing, and synthesizing information. As Sweller (1988) noted, complexity increases with greater cognitive resource demands, engaging working memory in reasoning and problem-solving. Similarly, Anderson et al. (2014) describes cognitive abilities as a continuum, from basic recall to higher-order thinking, with difficulty rising for tasks requiring analysis, synthesis, and evaluation. By mapping LLM prompting strategies onto this hierarchy, we systematically assess how LLMs handle varying cognitive loads. This framework provides a structured, cognitively grounded method for evaluating model performance across tasks of differing complexity. This study is directed by the following research questions:
Research Questions: RQ1: Does cognitively inspired strategic selection of prompts enable small language models (SLMs) to match the performance of LLMs? RQ2: How can cognitive demand measurements of LLMs guide model selection and deployment decisions beyond traditional metrics? RQ3: How can we align prompt complexity with task demands to optimize both computational efficiency and performance?
@@
This paper introduces the HPT, a set of rules that maps the human cognitive principles for assessing the complexity of different prompting strategies. It employs the HPF shown in Figure 1, a prompt selection framework that selects the prompt with the optimal cognitive load on LLM required in solving the task. The main contributions of this work are:
@@
  • Hierarchical Prompting Taxonomy (HPT): The paper introduces HPT, rules mapping prompting strategies to human cognitive principles, enabling a universal measure of LLMs’ task complexity.
  • Hierarchical Prompting Framework (HPF): The HPF framework selects the best prompt from five strategies to optimize LLMs’ cognitive load, improving evaluation and performance transparency.
@@
  • Hierarchical Prompting Index (HPI): HPI [^2] HPI can be quantitatively assessed to analyze the cognitive abilities of an LLM and the cognitive demands imposed by datasets on LLMs, as both factors are interchangeably related to the complexity of tasks. quantitatively assesses LLMs’ task complexity across datasets, revealing cognitive demands on various LLMs.
@@
Figure 2. Analogical framework comparing the HPF with "open book" examination methodology. The diagram illustrates how HPF components (below) mirror traditional educational assessment elements (above), with parallel relationships between task complexity levels, resource utilization (prompts/textbooks), and performance metrics (HPI/student effort). This comparison demonstrates how LLM task complexity scales similarly to educational assessment complexity, from simple lookup tasks to complex synthesis problems.
HPF can be compared to an "open book" exam (see Figure 2), with tasks analogous to questions and prompting strategies akin to textbooks. The exam questions, ranging from basic recall to complex analysis, parallel the cognitive challenges in HPT tasks. Similarly, textbooks offer structured support, much like HPF, which arranges prompts by complexity to assist LLMs. A glossary lookup represents a task with low complexity, whereas solving a multi-step analytical problem indicates high complexity. The effort exerted by a student is similar to HPI, which measures the cognitive demand on LLMs. Just as structured learning materials improve students’ performance, carefully crafted hierarchical prompts help LLMs in addressing increasingly complex tasks more effectively.
@@
The remainder of the paper is structured as follows: Section 2 reviews the related work on prompting and evaluation in LLMs. Section 3 details the HPT and its associated frameworks. Section 4 outlines the experimental setup, results, and ablation studies. Section 5 concludes the paper. Section 6 discusses the limitations of the work. Section 7 discusses the ethical impact of the work.
2. Related Work
@@ 2.1. Prompt Engineering
Prompting plays a vital role in unlocking the full potential of LLMs. By designing specific input prompts, the LLM’s responses can be guided, significantly influencing the quality and relevance of the output. Effective prompting strategies have enhanced LLM performance on tasks ranging from simple question-answering to complex reasoning and problem-solving. Recent research has explored various approaches to prompting and reasoning evaluation in LLMs. Chain-of-Thought (CoT) prompting (Wei et al., 2022b) elicits step-by-step reasoning, improving performance on complex tasks. Specializing smaller models (Fu et al., 2023) and using large models as reasoning teachers (Ho et al., 2022) have demonstrated the potential for enhancing reasoning capabilities. Emergent abilities in LLMs, which appear suddenly at certain scale thresholds, have also been a topic of interest. Wei et al. (2022a) examined these abilities in few-shot prompting, discussing the underlying factors and implications for future scaling. Complementing this, Kojima et al. (2022) demonstrated that LLMs could exhibit multi-step reasoning capabilities in a zero-shot setting by simply modifying the prompt structure, thus highlighting their potential as general reasoning engines. Yao et al. (2023) introduced the Tree-of-Thoughts framework, enabling LLMs to deliberate over coherent text units and perform heuristic searches for complex reasoning tasks. This approach generalizes over chain-of-thought prompting and has shown significant performance improvements in tasks requiring planning and search, such as creative writing and problem-solving games. Kong et al. (2024) introduced role-play prompting to improve zero-shot reasoning by constructing role-immersion interactions, which implicitly trigger chain-of-thought processes and enhance performance across diverse reasoning benchmarks. Progressive-hint prompting (Zheng et al., 2023) has been proposed to conceptualize answer generation and guide LLMs toward correct responses. Metacognitive prompting (Wang and Zhao, 2024) incorporates self-aware evaluations to enhance understanding abilities.
These studies highlight progress in using innovative prompting techniques to improve LLMs’ emergent abilities, reasoning, interaction strategies, robustness, and evaluation. Yet, challenges persist in prompt design, managing complex reasoning tasks, and performance evaluation across various scenarios. Although LLMs show promising emergent abilities, they frequently lack predictability and control, and their resistance to misleading prompts is still an issue.
2.2. Prompt Optimization and Selection
@@
The challenge of optimizing prompts for LLMs has been addressed in several key studies, each contributing unique methodologies to enhance model performance and efficiency. Shen et al. (2023) introduce PFLAT, a metric utilizing flatness regularization to quantify prompt utility, which leads to improved results in classification tasks. Do et al. (2024) propose a structured three-step methodology that contains data clustering, prompt generation, and evaluation, effectively balancing generality and specificity in prompt selection. ProTeGi (Pryzant et al., 2023) offers a non-parametric approach inspired by gradient descent, leveraging natural language "gradients" to iteratively refine prompts. Wang et al. (2024) present PromISe, which transforms prompt optimization into an explicit chain of thought, employing self-introspection and refinement techniques. Zhou et al. (2023a) proposed DYNAICL, a framework for efficient prompting that dynamically allocates in-context examples based on a meta-controller’s predictions, achieving better performance-efficiency trade-offs compared to uniform example allocation.
@@
These studies seek to automate prompt design, reducing reliance on manual trial-and-error while improving efficiency and scalability across tasks and models. They report performance gains of 5% to 31% across benchmarks, highlighting the growing significance of prompt optimization. Future research directions include exploring theoretical foundations, combining optimization techniques, and differentiating task-specific from general-purpose strategies.
2.3. Evaluation Benchmarks
@@
To facilitate the evaluation and understanding of LLM capabilities, Zhu et al. (2024) introduced PromptBench, a unified library encompassing a variety of LLMs, datasets, evaluation protocols, and adversarial prompt attacks. This modular and extensible tool aims to support collaborative research and advance the comprehension of LLM strengths and weaknesses. Further exploring reasoning capabilities, Qiao et al. (2023) categorized various prompting methods and evaluated their effectiveness across different model scales and reasoning tasks, identifying key open questions for achieving robust and generalizable reasoning. (Wang et al., 2021) introduced a multitask benchmark for LLM robustness evaluation, which extends the original GLUE (Wang et al., 2018) benchmark to assess model robustness against adversarial inputs. It incorporates perturbed versions of existing GLUE tasks, such as paraphrasing, negation, and noise, to test models’ abilities with challenging data. The study highlights that despite their success on clean datasets, state-of-the-art models often struggle with adversarial examples, underscoring the importance of robustness evaluations in model development.
3. Hierarchical Prompting Taxonomy
@@
Figure 3. Hierarchical Prompting Taxonomy: A taxonomy designed to assess the complexity of prompting strategies based on the criteria: Basic Recall and Reproduction, Understanding and Interpretation, Analysis and Reasoning, and Application of Knowledge and Reasoning.
3.1. Governing Rules
@@
  1. (1)
    Basic Recall and Reproduction: This reflects the fundamental cognitive process of remembering and reproducing factual information without analysis or interpretation, which involves mere recognition or retrieval of knowledge from memory (Anderson et al., 2014).
  2. (2)
    Understanding and Interpretation: This corresponds to the second cognitive rule of (Bloom, 1956), where individuals must not only recall information but also explain it in their own words, summarize key points or clarify the meaning of content. This rule demands an intermediate cognitive load involving information processing rather than retrieving it.
  3. (3)
    Analysis and Reasoning: This aligns with the analysis stage of (Bloom, 1956), which involves higher cognitive functions such as comparison, contrast, and deep understanding of the underlying principles. It is more complex than mere understanding because it requires examining structure and identifying patterns and connections.
  4. (4)
    Application of Knowledge and Execution: This mirrors the application and evaluation stages of (Bloom, 1956), where individuals must not only understand and analyze but also use knowledge to perform multi-step tasks, solve complex problems, and execute decisions. It represents the most cognitively complex tasks, which require synthesis of information and practical decision-making, highlighting the critical leap from understanding theory to executing it in practice.
@@
3.2. Hierarchical Prompting Framework
@@
The five prompting strategies were selected to ensure comprehensive coverage of cognitive demands rather than maximizing the number of strategies (see Appendix A). This makes HPF adaptable, allowing for replication or expansion with similar strategies. The strategies, ordered by increasing complexity, are:
  1. (1)
    Role Prompting (Kong et al., 2024): Specifies the LLM’s role in task resolution, exerting minimal influence from HPT principles.
  2. (2)
    Zero-Shot Chain-of-Thought Prompting (Zero-CoT) (Kojima et al., 2022): Uses “Let’s think step by step” to encourage reasoning, moderately influenced by rule 3.
  3. (3)
    Three-Shot Chain-of-Thought Prompting (3-CoT) (Wei et al., 2022b): Provides three examples to guide reasoning, strongly influenced by rules 1 and 2, with moderate influence from rule 3.
  4. (4)
    Least-to-Most Prompting (Zhou et al., 2023b): Breaks tasks into sub-problems, requiring recall, interpretation, and analysis, exerting strong influence from rules 1, 2, & 3.
  5. (5)
    Generated Knowledge Prompting (GKP) (Liu et al., 2022): Integrates external knowledge, demanding correlation, application, and analysis, making it the most cognitively complex (rules 2, 3, and 4). Llama-3 8B generates the external knowledge in experiments.
3.3. Hierarchical Prompting Index
HPI is an evaluation metric for assessing the task complexity of LLMs over different datasets, which is influenced by the HPT rules. A lower HPI for a dataset suggests that the corresponding LLM is more adept at solving the task with fewer cognitive processes. For each dataset instance, we begin with the least complex prompting strategy and progressively move through the HPF prompting strategies until the instance is resolved. The HPI corresponds to the complexity level of the prompting strategy where the LLM first tackles the instance.
Algorithm 1 HPI Computation